# FedSDAF: Leveraging Source Domain Awareness for Enhanced Federated Domain Generalization

### v3

@@ B.1 Fine-grained Ablation Study on Hard Domains
The results are presented in Table 5. On these hard domains, the contribution of each component becomes significantly more pronounced. For instance, enabling BKD and MHSA leads to substantial gains of 3-7% on VLCS and OfficeHome. The full model consistently achieves the best results, with the MHSA mechanism being particularly crucial for boosting performance on these challenging generalization tasks. This underscores the importance of each component in our framework for achieving robust generalization, an effect that is most visible where it matters most.
Table 5: Fine-grained ablation study on the most challenging domains. The performance improvements for each component are more significant on these domains compared to the overall average.
@@
BKD MHSA $\boldsymbol{\hat{A}_{di}}$ $\boldsymbol{\hat{A}_{dw}}$ PACS (P) VLCS (L) OfficeHome (C)
@@ B.2 Visualization
We present several visualization results to provide an intuitive understanding of the advantages of FedSDAF. In Fig. 6, we apply the t-SNE algorithm to visualize the features of target samples extracted by ViT-B/16, Domain Invariant Adapter, Domain Aware Adapter, and FedSDAF (KIA), respectively, with Sketch designated as the target domain on PACS. The t-SNE visualization reveals critical insights into FedSDAF’s feature discrimination capability within the challenging Sketch domain. While ViT-B/16 (Fig.6d) exhibits dispersed feature distribution with substantial inter-class overlap, both individual adapters demonstrate distinct adaptation strategies aligned with our design motivation: The Domain Invariant Adapter (Fig.6a) produces elliptical feature clusters with overlapping boundaries, reflecting its emphasis on cross-domain alignment at the expense of category discrimination. Conversely, the Domain Aware Adapter (Fig.6b) generates compact but partially misaligned clusters, indicative of its focus on capturing source-specific discriminative features while neglecting domain shift mitigation. FedSDAF (Fig.6c) fundamentally overcomes these limitations through hierarchical feature disentanglement. The isolated red cluster in the lower-left quadrant manifests the domain-aware component’s success in preserving source-discriminative patterns, while the uniform spacing between blue and purple clusters in the upper region verifies the domain-invariant module’s effectiveness in establishing cross-domain consistency. This dual mechanism achieves geometric synergy - the domain-aware adapter actively pushes class clusters apart by amplifying source-specific discriminators, while the domain-invariant adapter passively pulls features toward domain-agnostic decision boundaries. Such collaborative dynamics explain why FedSDAF surpasses both the base model and standalone adapters, as visualized by the sharpened cluster margins and reduced overlap density compared to other configurations. These visual patterns collectively validate our core motivation: Source-aware feature refinement requires simultaneous discriminative pattern preservation (through domain-aware adaptation) and cross-domain topology alignment (via invariant feature projection), which FedSDAF achieves through its dual-branch architecture.
B.3 Scalability with Increasing Client Pool Size
@@
The results, presented in Figure 7, demonstrate that FedSDAF exhibits robust scalability. As the number of clients increases, there is a slight and graceful degradation in accuracy across all domains. This trend is expected in federated settings, as a larger client pool typically introduces greater data heterogeneity and makes achieving a global consensus more challenging. However, the performance decline is modest, even when scaling to 50 clients. For instance, on the most challenging Sketch (S) domain, the accuracy remains high, showcasing the framework’s ability to maintain its effectiveness. This highlights that FedSDAF’s core mechanisms, particularly the synergistic learning between the domain-aware and domain-invariant adapters, are resilient to increased decentralization, making the framework highly suitable for large-scale, real-world deployments.
@@
B.4 Analysis of Adapter Fusion Strategy
Our framework integrates the outputs of the Domain-Aware ($A_{dw}$) and Domain-Invariant ($A_{di}$) adapters using a simple summation with equal weights, as described in Equation (2): $h^{\prime}=h+\frac{1}{2}\hat{A}_{dw}(h)+\frac{1}{2}\hat{A}_{di}(h)$. To validate this design choice, we performed an ablation study to analyze the impact of different fusion strategies on the PACS dataset. We explored two alternatives: (1) modifying the relative weighting of the adapter outputs, and (2) replacing the summation with a concatenation operation.
@@ B.4 Analysis of Adapter Fusion Strategy
B.5 Ablation Study on MHSA via Dynamic Distillation Weighting
In our main experiments, the FedSDAF framework utilizes a Multi-Head Self-Attention (MHSA) mechanism alongside a fixed weight for the bidirectional knowledge distillation (BKD) loss. To validate the unique contribution of the MHSA component, we conducted an ablation study where we removed it and attempted to compensate by introducing a more sophisticated, dynamic weighting strategy for the distillation loss. Specifically, instead of a fixed $\alpha$, we employed an exponentially increasing weight governed by the formula:
@@ B.5 Ablation Study on MHSA via Dynamic Distillation Weighting
The analysis of this experiment leads to a clear conclusion. The heatmap reveals that the optimal performance for the MHSA-free variant is achieved with a high $\alpha_{\text{max}}$ (2.0) and a moderately fast rate of increase ($\tau=100$), yielding a peak average accuracy of approximately 97.08%. While this result confirms that a carefully tuned dynamic distillation schedule can enhance performance, it crucially represents the performance ceiling for this ablated model configuration. Notably, this peak accuracy remains significantly below the performance of our full FedSDAF model, which incorporates the MHSA mechanism (as detailed in Table 5). This performance deficit underscores that the contribution of MHSA is not merely a matter of balancing feature fusion, which could be emulated by adjusting loss weights. Instead, it provides a distinct architectural advantage, likely by capturing complex inter-feature relationships that are essential for robust generalization. Therefore, this ablation study validates that while a dynamic distillation weight is a viable optimization strategy, it is not a substitute for the fundamental benefits conferred by the MHSA component, thereby justifying its integral role in our proposed framework.

### v4

@@ IV-H1 Fine-grained Ablation Study on Hard Domains
The results are presented in Table V. On these hard domains, the contribution of each component becomes significantly more pronounced. For instance, enabling BKD and MHSA leads to substantial gains of 3–7% on VLCS and OfficeHome. The full model consistently achieves the best results, with the MHSA mechanism being particularly crucial for boosting performance on these challenging generalization tasks. This underscores the importance of each component in our framework for achieving robust generalization, an effect that is most visible where it matters most.
TABLE V: Fine-grained ablation study on the most challenging domains. The performance improvements for each component are more significant on these domains compared to the overall average.
@@
BKD MHSA $\boldsymbol{A_{di}}$ $\boldsymbol{A_{dw}}$ PACS (P) VLCS (L) OfficeHome (C)
@@ IV-H2 Visualization
We present several visualization results to provide an intuitive understanding of the advantages of FedSDAF. In Fig. 7, we apply the t-SNE algorithm to visualize the features of target samples extracted by ViT-B/16, the Domain-Invariant Adapter, the Domain-Aware Adapter, and FedSDAF (KIA), respectively, with Sketch designated as the target domain on PACS. The t-SNE visualization reveals critical insights into FedSDAF’s feature discrimination capability within the challenging Sketch domain. While ViT-B/16 (Fig. 7(d)) exhibits a dispersed feature distribution with substantial inter-class overlap, both individual adapters demonstrate distinct adaptation strategies aligned with our design motivation. The domain-invariant adapter $\hat{A}_{di}$ (Fig. 7(a)) produces elliptical feature clusters with overlapping boundaries, reflecting its emphasis on cross-domain alignment at the expense of category discrimination. Conversely, the domain-aware adapter $\hat{A}_{dw}$ (Fig. 7(b)) generates compact but partially misaligned clusters, indicative of its focus on capturing source-specific discriminative features while neglecting domain shift mitigation. FedSDAF (Fig. 7(c)) fundamentally overcomes these limitations through hierarchical feature disentanglement. The isolated red cluster in the lower-left quadrant manifests the domain-aware component’s success in preserving source-discriminative patterns, while the uniform spacing between blue and purple clusters in the upper region verifies the domain-invariant module’s effectiveness in establishing cross-domain consistency. This dual mechanism achieves geometric synergy: the domain-aware adapter $\hat{A}_{dw}$ actively pushes class clusters apart by amplifying source-specific discriminators, while the domain-invariant adapter $\hat{A}_{di}$ passively pulls features toward domain-agnostic decision boundaries. Such collaborative dynamics explain why FedSDAF surpasses both the base model and standalone adapters, as visualized by the sharpened cluster margins and reduced overlap density compared to other configurations. These visual patterns collectively validate our core motivation: source-aware feature refinement requires simultaneous discriminative pattern preservation (through domain-aware adaptation) and cross-domain topology alignment (via invariant feature projection), which FedSDAF achieves through its dual-branch architecture.
IV-H3 Scalability with Increasing Client Pool Size
@@
The results, presented in Figure 8, demonstrate that FedSDAF exhibits robust scalability. As the number of clients increases, there is a slight and graceful degradation in accuracy across all domains. This trend is expected in federated settings, as a larger client pool typically introduces greater data heterogeneity and makes achieving a global consensus more challenging. However, the performance decline is modest, even when scaling to 50 clients. For instance, on the most challenging Sketch domain, the accuracy remains high, showcasing the framework’s ability to maintain its effectiveness. This highlights that FedSDAF’s core mechanisms, particularly the synergistic learning between the domain-aware and domain-invariant adapters, are resilient to increased decentralization, making the framework highly suitable for large-scale, real-world deployments.
@@
IV-H4 Analysis of Adapter Fusion Strategy
Our framework integrates the outputs of the domain-aware ($A_{dw}$) and domain-invariant ($A_{di}$) adapters using a simple summation with equal weights, as described in Equation (2): $h^{\prime}=h+\frac{1}{2}\hat{A}_{dw}(h)+\frac{1}{2}\hat{A}_{di}(h)$. To validate this design choice, we performed an ablation study to analyze the impact of different fusion strategies on the PACS dataset. We explored two alternatives: modifying the relative weighting of the adapter outputs, and replacing the summation with a concatenation operation.
@@ IV-H4 Analysis of Adapter Fusion Strategy
IV-H5 Ablation Study on MHSA via Dynamic Distillation Weighting
In our main experiments, the FedSDAF framework utilizes a MHSA mechanism alongside a fixed weight for the BKD loss. To validate the unique contribution of the MHSA component, we conducted an ablation study where we removed it and attempted to compensate by introducing a more sophisticated, dynamic weighting strategy for the distillation loss. Specifically, instead of a fixed $\alpha$, we employed an exponentially increasing weight governed by the formula:
@@ IV-H5 Ablation Study on MHSA via Dynamic Distillation Weighting
The analysis of this experiment leads to a clear conclusion. The heatmap reveals that the optimal performance for the MHSA-free variant is achieved with a high $\alpha_{\text{max}}$ (2.0) and a moderately fast rate of increase ($\tau=100$), yielding a peak average accuracy of approximately 97.08%. While this result confirms that a carefully tuned dynamic distillation schedule can enhance performance, it crucially represents the performance ceiling for this ablated model configuration. Notably, this peak accuracy remains significantly below the performance of our full FedSDAF model, which incorporates the MHSA mechanism (as detailed in Table V). This performance deficit underscores that the contribution of MHSA is not merely a matter of balancing feature fusion, which could be emulated by adjusting loss weights. Instead, it provides a distinct architectural advantage, likely by capturing complex inter-feature relationships that are essential for robust generalization. Therefore, this ablation study validates that while a dynamic distillation weight is a viable optimization strategy, it is not a substitute for the fundamental benefits conferred by the MHSA component, thereby justifying its integral role in our proposed framework.
V Conclusion